OpenStudy (anonymous):

Solve 3x^2 = -30x + 24

OpenStudy (abb0t):

add 30\(x\) and subtract \(24\) from both sides. set it equal to zero. solve for the polynomial

OpenStudy (abb0t):

\(3x^2+30x+24=0\) and it looks like you can factor out a 3 to get the first term: \(x^2\) by itself to make it easier for you.

OpenStudy (anonymous):

1. -10+or-sqrt33/2 2. -5+or-2sqrt33 3.-10+or-sqrt33 4.-5+or-sqrt33

OpenStudy (abb0t):

I already told you what to do. now, just use the quadratic formula to solve: \(\large \frac{-b \pm \sqrt{b^2-4ac}}{2a}\) where you have \(ax^2+bx+c\)

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

the are the answers i got

OpenStudy (abb0t):

Plug it in. It's not that hard, daryl.

OpenStudy (anonymous):

i did whats the answer?

OpenStudy (abb0t):

You tell me.